# Query Rewriting via LLMs

### v1

@@ Abstract
Query rewriting is a classical technique for transforming complex declarative SQL queries into “lean” equivalents that are conducive to (a) faster execution from a performance perspective, and (b) better understanding from a developer perspective. The rewriting is typically achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system. In recent times, LLM-based techniques have also been mooted, but they are prone to both semantic and syntactic errors.
We investigate here, how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency. Our study shows that these goals can be progressively achieved through incorporation of (a) an ensemble suite of basic prompts, (b) database-sensitive prompts via redundancy removal and selectivity-based rewriting rules, and (c) LLM token probability-guided rewrite paths. Further, a suite of statistical and logic-based tools can be used to guard against errors produced by the model.
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from multiple benchmarks on contemporary database platforms. The results show significant improvements over SOTA rewriting techniques – for instance, on TPC-DS, LITHE constructed productive ($>$1.5x speedup) rewrites for two-thirds of the query suite, delivering four times more coverage than SOTA. Further, the geometric mean of its estimated execution speedups was an order-of-magnitude jump over SOTA performance. In essence, LITHE offers a potent and robust LLM-based intermediary between enterprise applications and database engines.
@@ 1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools. A compelling case in point is the blog-processing query [4] shown in Figure 2 – this query was produced by the popular Entity Framework [21] mapper. It computes a daily summary of rating metrics for highly-rated blogs. This seemingly complex nested query can be equivalently rewritten in the “lean” flat version shown in Figure 2.
@@
FROM Posts AS p0 INNER JOIN Blogs AS b1 ON
p0.BlogId = b1.BlogId
@@
Albeit less common, we observe such superfluity even in the hand-crafted expert query domain – for instance, TPC-DS Query Q23 includes a sub-query that identifies customers with the maximum store sales in a four-year period; this sub-query has a redundant join of store_sales with customer despite the FK-PK relationship between these tables, as highlighted in [25].
Transforming bloated queries to lean equivalents in the application source code is beneficial from multiple perspectives: First, rewrites can improve code readability, a major concern in industrial contexts. Second, query optimizers are expected to, in principle, remove bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor query performance. In fact, both examples discussed above were implemented as such by the current PostgreSQL optimizer (v16) [1], without any change. Therefore, a query rewriter can serve as an effective and non-invasive mechanism for delivering good performance despite inherent optimizer limitations.
Given this motivational context, a viable SQL-to-SQL query transformer should satisfy the following criteria: (1) The transformed query should be semantically equivalent to the original; (2) There should ideally be performance improvement due to the rewrite, but at the very least, no regression in query performance; and (3) The transformation overheads must be practical for deployment.
@@
A variety of innovative Rule-based (e.g. Learned Rewrite [38]) and Model-based (e.g. Gen-Rewrite [19]) approaches have been proposed in the literature for SQL query transformations. However, as explained later in the related work section (Section 8), the benefits of these state-of-the-art (SOTA) techniques are curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space.
An alternative and extremely simple strategy would be to take advantage of the large body of recent literature on LLM-based Text-to-SQL translations (e.g. [17, 35, 33]). That is, first invoke an LLM to transform the original SQL to text, and then employ a Text-to-SQL tool to produce quality SQL. However, as discussed later in Section 7, we found this approach to be highly prone to semantic errors and lacking execution benefits. This is perhaps to be expected given the focus of Text-to-SQL tools on query formulation, orthogonal to our objective of performance enhancement. Therefore, the SQL-to-SQL problem merits a fresh assessment from first principles.
@@
In this study, we carry out a calibrated investigation of how LLMs could be reliably used to produce performant rewrites of complex SQL queries. The proposed techniques, described below, have been implemented and evaluated in the LITHE (LLM Infused Transformations of Hefty queries) rewriting system developed by our group.
Basic Prompts. We begin with an ensemble suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs. Interestingly, we find that more information is not necessarily better wrt rewriting quality, and the best prompt granularity is query-specific. Moreover, this simple prompt ensemble was found to deliver performance similar to the SOTA techniques.
Database-sensitive Prompts. To help the model adapt to different query patterns and structures, we next introduce rules in the prompts. But our choice and usage of rules are markedly different from those used in prior work. First, our rules are invoked directly in query space. The LLM therefore gains the latitude to generalize the rule usage to a wide range of queries. This is in contrast to the hardwired and narrow rule application mechanisms (e.g. Calcite [8] rules) typically used in existing rewrite systems, which operate in plan space. Second, to minimize rule application overheads, we incorporate only a handful of rules that serve to eliminate the primary causes of poor query performance. Whereas current rewrite systems employ a large number of rules in their processing. In particular, we work with two classes of rules:
@@
2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
@@
Our experiments are carried out on a variety of synthetic and real benchmarks: TPC-DS [11], DSB [12], JOB [15], StackOverflow [20] and Archer [37], covering a broad spectrum of databases and queries. Further, anonymized versions, where the table and column names are altered to not convey any semantic meaning about their contents, are also evaluated to assess robustness to databases not seen by the LLM during its training phase. These databases are hosted on PostgreSQL and GPT-4o is used as the LLM.
We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [38], LLM-R2 [18], GenRewrite [19]) as well as a baseline LLM prompt [19]. The primary metrics are the execution cost reductions and the rewriting overheads. For LLM-based techniques, the number of tokens in the prompts is also monitored since the financial charges for LLM usage are typically dependent on this number. To understand the impact of the various components of LITHE, a systematic ablation study is carried out. Finally, the performance on alternative database and LLM platforms is also evaluated.
Verifying Semantic Equivalence. The general problem of proving query equivalence has long been known to be computationally hard [6]. However, for restricted classes of queries, various logic-based equivalence-proving tools are available (e.g. Cosette [10], SQL-Solver [13], and QED [30]). In our LITHE system, we use QED [30] to verify rewrite correctness for queries within its scope. The remainder are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases using small sampled versions of the original database. We also provide the user with an option to test result equivalence using the original database.
Results. Our experiments demonstrate that LITHE achieves semantically correct transformations that deliver a significant reduction in estimated execution times for many queries. Moreover, LITHE performed better than or at least as well as SOTA on all the benchmarks. As a compelling case in point, LITHE constructed “productive” (¿1.5x speedup) rewrites for about two-thirds of the complete TPC-DS query suite, delivering four times more coverage than SOTA techniques. Further, the GM (Geometric Mean) of the reduction in query execution costs was 30.6, an order-of-magnitude jump over that offered by SOTA.
In essence, LITHE offers a potent LLM-based intermediary between enterprise applications and database engines.
@@
    Assesses LLM suitability for SQL-to-SQL transformation.
@@
    Transforms directly in query space instead of a plan space intermediate representation, leading to robust rewrites.
@@
    Incorporates potent database-sensitive rules in LLM prompts covering schematic and statistical dimensions.
@@
    Evaluates rewriting robustness over a broad range of database environments and benchmarks.
@@
    Demonstrates substantive performance benefits, well beyond SOTA, from LLM-powered query rewrites.
@@ 2 LITHE Overview
The LITHE architecture, illustrated in Figure 3, consists of two main components: A prompt-based rewriting pipeline, and a token probability-driven rewrite exploration.
Prompt-based rewriting pipeline. This pipeline comprises four modules: (1) LLM Prompting, (2) Syntax Verification, (3) Query Costing, and (4) Semantic Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output is checked by the syntax verification component – if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively until a valid SQL query is obtained. Then, the execution cost of this version is evaluated – if more than the cost of the original query, the rewrite is discarded. Otherwise, the rewrite is tested for semantic equivalence. If a match is established, the rewrite is returned; otherwise, the original query itself is returned. Finally, to prevent runaway situations, a threshold is set on the number of syntax correction attempts (5 in our experiments) and the original query is returned if this threshold is breached.
The above process is repeated for each of the different prompts described in Sections 3 and 4. Then, the prompt providing the least-cost rewrite is returned. Since each prompt is evaluated within a fresh LLM context, there is no dependency on the order in which they are assessed.
Within the pipeline, LITHE uses (a) the query parser from the database engine for syntax verification, (b) the query optimizer for cost estimation, and (c) the techniques described in Section 6.3 to check semantic equivalence.
Figure 3: High-level architecture of LITHE
Token probability-driven rewrite. The pipeline-chosen prompt is used as an input to a Monte-Carlo tree search (MCTS)-based procedure to further improve the query rewrite quality. Specifically, multiple exploration paths are followed whenever the LLM lacks high confidence in the predicted token (details in Section 5).
@@ 2.1 Performance Framework
As mentioned in the Introduction, the majority of our experiments are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine.
We define a productive rewrite (PR) as one that improves the query’s performance by at least 1.5 times wrt the execution times estimated by the database engine’s query optimizer – this aggressive choice of threshold is to make the rewrite outcome desirable from an industrial perspective. Given a query workload, the set of feasible productive rewrites (FPR) consists of the queries for which PR are known to be feasible by some technique or the other. Note that it may be inherently impossible to produce PR on queries that are structurally simple or efficiently written by human experts. Finally, a maximum productive rewrite (MPR) is when the PR speedup of the rewrite matches the best known PR on this query.
The overall benefit provided by a rewriting tool is quantified by the number of PR obtained on the query workload. Additionally, we also measure SpeedupGM, the geometric mean (GM) of the speedups obtained on the FPR queries in this workload. (The geometric mean is used to ensure equal weight for all queries in the workload, irrespective of their relative durations.)
@@ 2.2 Query Micro-benchmark
To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft productive rewrites. These human rewrites deliver a SpeedupGM of 13.9 on the micro-benchmark, serving as an informal “upper-bound” on what is computationally attainable. Later, in Section 7, we extend the evaluation to a suite of complete benchmarks.
@@ 3 Basic Prompts
In this section, we explore the simplest interface to LLMs, namely prompting, for query rewriting. We evaluate four basic prompts, illustrated in Figure 4, which cover a progressive range of detail in the instructions and test the effectiveness of the LLM’s base knowledge.
@@
Prompt 3: Verbose instructions are given to rewrite the query, providing step-by-step guidance to the LLM to think rationally. It is first asked to pick out potential inefficiencies in the input query, and then tasked to identify approaches to address these inefficiencies. Finally, it is instructed to apply the identified solution. Essentially, the prompt tries to make the LLM reason akin to a human expert.
@@
The performance of the four prompt templates on the micro-benchmark (where all 10 queries are in FPR) is shown in Table 1. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that each prompt does the best on certain queries – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. While this ensemble approach (Row 5 in Table 1) raises the PR to 6, there remain four queries that could not be productively rewritten by the prompts alone.
Table 1: Performance of Basic Prompts
Prompt # PR SpeedupGM
Prompt 1 3 2.7
Prompt 2 4 1.57
Prompt 3 3 1.37
Prompt 4 2 1.27
Prompt Ensemble 6 3.6
SOTA Ensemble 5 3.4
The SpeedupGM, shown in the last column, for each individual prompt is less than 3, while that of the ensemble reaches 3.6. But these speedups, while productive, are all lower than those of the human rewrites, that is, they are not MPR.
Finally, as another baseline, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform. They delivered (Row 6 in Table 1) 5 productive rewrites with a GM of 3.4, indicating the wide gap between the current state-of-the-art and what is possible.

### v2

@@ Abstract
When complex SQL queries suffer slow executions despite query optimization, DBAs typically invoke automated query rewriting tools to recommend “lean” equivalents that are conducive to faster execution. The rewritings are usually achieved via transformation rules, but these rules are limited in scope and difficult to update in a production system. Recently, LLM-based techniques have also been suggested, but they are prone to semantic and syntactic errors.
We investigate here how the remarkable cognitive capabilities of LLMs can be leveraged for performant query rewriting while incorporating safeguards and optimizations to ensure correctness and efficiency. Our study shows that these goals can be progressively achieved through incorporation of (a) an ensemble suite of basic prompts, (b) database-sensitive prompts via redundancy removal and selectivity-based rewriting rules, and (c) LLM token probability-guided rewrite paths. Further, a suite of logic-based and statistical tools can be used to check for semantic violations in the rewrites prior to DBA consideration.
We have implemented the above LLM-infused techniques in the LITHE system, and evaluated complex analytic queries from standard benchmarks on contemporary database platforms. The results show significant performance improvements for slow queries, with regard to both abstract costing and actual execution, over both SOTA techniques and the native query optimizer. For instance, with TPC-DS on PostgreSQL, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison.
Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise query performance.
@@ 1 Introduction
The SQL queries embedded in enterprise applications are often riddled with inefficiencies and redundancies, especially when machine-generated by modeling software such as ORM tools. Query optimizers are expected to, in principle, automatically remove such wasteful bloat while constructing efficient execution plans. However, in practice, they are often led astray by complex representations, resulting in poor response times.
@@
FROM Posts AS p0 INNER JOIN Blogs AS b1
ON p0.BlogId = b1.BlogId
@@
As a case in point, consider the blog-processing query shown in Figure 2, which computes a daily summary of rating metrics for highly-rated blogs. When this seemingly complex query is given to the current PostgreSQL engine (v16) [1], its execution plan essentially mimics the hierarchical structure of the query. This strategy leads to multiple scans and joins of the Blogs and Posts tables, making the query take several minutes to complete. However, the query can be equivalently rewritten (assuming NOT NULL column constraints) in the “lean” flat version shown in Figure 2 – this alternative requires only a single join, and runs more than an order-of-magnitude faster, completing within seconds!
For such optimizer-failure scenarios, an alternative option to rectify slow-running queries is to carry out external tuning – in particular, DBAs typically resort to SQL-to-SQL (S2S) query rewriting tools in the hope that they may recommend performant alternatives. For this context, a viable S2S query transformer should satisfy the following criteria: (1) A wide range of queries should be covered in its scope; (2) The transformed query should be semantically equivalent to the original; (3) The rewrite should ideally improve performance, but at least not cause regression; and (4) The transformation overheads must be practical for deployment.
@@
A variety of innovative Rule-based (e.g. Learned Rewrite [45]) and Model-based (e.g. Gen-Rewrite [25]) approaches have been proposed in the literature for S2S transformations. Further, a recent vision paper [46] has persuasively advocated the use of LLMs for such query rewriting. These novel state-of-the-art (SOTA) techniques have foregrounded the potential benefits of query rewriting. However, as explained later in Section 9, their realization of these benefits is curtailed by: (a) restrictions in rewrite scope, (b) susceptibility to semantic and syntactic errors, and (c) transformations via the plan space rather than directly in query space – the latter provides greater scope for candidate rewrites since only intent, and not implementation, is expressed. In this paper, we address these lacuna and amplify the effectiveness of query rewriting.
Figure 3: High-level architecture of LITHE
@@
We propose LITHE (LLM Infused Transformations of HEfty queries), an LLM-based query rewriting assistant to aid DBAs in tuning slow-running queries that have entailed their intervention. As illustrated in the architectural diagram of Figure 3, LITHE takes as input the user query $Q^{U}$ and outputs a transformed query $Q^{T}$, together with (a) the expected performance improvement, in terms of optimizer estimated cost, of $Q^{T}$; (b) a verification label indicating the mechanism (provable or statistical) used to determine that $Q^{T}$ is semantically equivalent to $Q^{U}$; and (c) a reasoning for why the LLM expects $Q^{T}$ to be helpful wrt performance. Armed with this information, DBAs can leverage their expertise to decide whether or not to use $Q^{T}$.
Our design of LITHE is based on a calibrated investigation of the suite of techniques described below:
Basic Prompts. We begin with a suite of generic prompts that cover a spectrum of detail, ranging from a single summary sentence to detailed instructions running to several paragraphs. Interestingly, we find that more information is not necessarily better wrt rewriting quality, and that the best prompt granularity is query-specific. Moreover, this basic prompt ensemble was found to itself deliver performance similar to the SOTA techniques.
Database-sensitive Prompts. To help the model adapt to different query patterns and structures, we next introduce rules in the prompts. Our rules are invoked directly in query space, providing the LLM with the latitude to generalize the rule usage to a wide range of queries. This is in contrast to the hardwired and narrow rule application mechanisms (e.g. Calcite [9] rules) typically used in existing rewrite systems, which operate in plan space.
In particular, we work with two classes of rules:
@@
2. Metadata-infused Rules: These rules make use of the rich metadata available in database environments, such as the logical schema (table definitions and constraints) and predicate selectivities, and include this information in the LLM prompts. To our knowledge, such metadata inclusion in prompts has not been considered before in the SQL rewrite context. As shown later in our experiments, it proves to be a powerful mechanism for ensuring performant rewrites across database environments.
@@
Semantic Equivalence Labels. After generating a rewrite, it is imperative to check for semantic equivalence to the original query. Unfortunately, the general problem of proving query equivalence is NP-complete [6]. However, for restricted classes of queries, various logic-based equivalence-proving tools are available, and in LITHE, we use the recently developed QED tool [36] to verify rewrite correctness for queries within its scope.
The remaining out-of-scope queries are checked, albeit not provably, by computing result equivalences on a diverse cluster of databases constructed using down-sampled versions of the original database. Finally, result equivalence can be optionally tested on the original database itself. While these checks do not suffer from false negatives, they may (rarely) incur false positives.
After carrying out the above exercise, LITHE provides a verification label – provable or statistical – regarding semantic equivalence. Ultimately, it is left to the DBA’s discretion to determine the validity and utility of the rewrite.
@@
Our first set of experiments to evaluate LITHE’s performance is carried out on the industry standard TPC-DS benchmark [12], hosted on the PostgreSQL platform with GPT-4o used as the LLM. The evaluation focuses on slow queries taking more than a threshold time to complete. We compare the performance of LITHE against SOTA techniques (specifically, Learned Rewrite [45], LLM-R2 [24], GenRewrite [25], as well as a baseline LLM prompt [25]). The primary metrics are (a) reductions in optimizer-estimated costs, (b) run-time speedups, and (c) rewriting overheads. For LLM-based techniques, the number of tokens used is also monitored since the financial charges for LLM usage are typically dependent on this number. To understand the independent utility of the various components of LITHE, a systematic ablation study is carried out.
Generalizability.
In our second stage of experiments, we evaluate generalizability of the above outcomes in a variety of new scenarios, including (a) commercial database engines, (b) unseen database schemas, and (c) alternative LLM platforms.
Results. Our experiments demonstrate that LITHE achieves, for many slow queries, semantically correct transformations that significantly reduce the abstract costs. In particular, for TPC-DS, LITHE constructed “highly productive” ($>1.5x$ estimated speedup) rewrites for as many as 26 queries, whereas SOTA promised such rewrites for only about half the number. Further, the GM (Geometric Mean) of LITHE’s cost reductions reached 11.5, almost double the 6.1 offered by SOTA.
We also evaluated whether the above cost reductions translated to real execution speedups. Here, we find that LITHE is indeed often substantively faster at run-time as well. Specifically, the geometric mean of the runtime speedups for slow queries was as high as 18.4 over the native optimizer, whereas SOTA delivered 6 in comparison. However, we also found that the well-known discrepancies between optimizer cost predictions and actual execution times [26] could be exacerbated by an injudicious choice of rule formulations. This separation suggests a future line of research to reduce the reasoning distance between language-based foundation models and statistics-based optimizer models.
Overall, LITHE is a promising step toward viable LLM-based advisory tools for ameliorating enterprise application performance.
@@
    Assesses LLM suitability for S2S transformation.
@@
    Transforms directly in query space instead of plan space intermediates, leading to performant rewrites.
@@
    Incorporates database-sensitive rules in LLM prompts, covering both schematic and statistical dimensions.
@@
    Evaluates rewrite quality over a broad range of database environments, demonstrating substantial benefits over both SOTA and the native optimizer.
@@
    Identifies learnings that could help guide research directions for industrial-strength query rewriting.
@@ 2 LITHE Overview
The LITHE architecture, illustrated in Figure 3, consists of the following five-stage pipeline:
1. Prompt-based rewriting. This component consists of two modules: LLM Prompting and Syntax Verification. The user query and an initial prompt are fed to the LLM prompting module, requesting a rewrite. The LLM output’s syntax is then verified with the database engine’s parser– if invalid, the error is fed back to the LLM via an updated prompt asking for a correction, and this goes on iteratively (within a threshold) until a valid SQL query is obtained.
This process is repeated for each of the different prompts described in Sections 3 and 4, and the resulting syntactically-valid rewrites are fed to the query costing component. We evaluate each prompt within a fresh LLM context, thereby making the prompt processing to be order-independent.
2. Query Costing. The costs of candidate rewrites are evaluated via the database engine’s optimizer. Rewrites whose costs are greater than that of the original query are immediately discarded. Whereas, the potentially beneficial rewrites (if any) are checked for semantic equivalence to the original query.
3. Semantic Equivalence. A suite of provable (logic-based) and statistical (result-based) techniques are employed to assess the semantic equivalence of a recommended rewrite. If the rewrite is deemed valid by this module, it is returned along with the prompt that generated it; otherwise an invalid label is returned.
4. Token probability-driven rewrite. The prompt producing the most beneficial (and valid) rewrite is used as input to a Monte-Carlo tree search (MCTS)-based procedure to further refine the rewrite quality. Specifically, multiple exploration paths are followed whenever the LLM lacks high confidence in the predicted token (details in Section 5). The query costing and semantic equivalence modules are also used internally within this procedure.
5. Output. The least expensive valid rewrite as identified by the MCTS module is returned to the DBA together with the cost of the rewritten query and the approach used to claim equivalence. Additionally, an LLM-generated reasoning for the expected performance improvement is also provided. If no valid rewrite is identified, the original query itself is returned.
@@ 2.1 Performance Framework
We consider a query that takes more than T seconds to complete on the native database engine as a “slow query”, potentially triggering intervention by the DBA. Based on common industry perception (e.g. [4]), T is set at 10 seconds in our study. For this context, we define a Cost Productive Rewrite (CPR) as a rewrite that improves a slow query’s performance by at least 1.5 times wrt the optimizer-estimated cost – this aggressive choice of threshold is so that: (a) there is enough headroom in the optimizer prediction that a runtime regression is unlikely; and (b) the benefits of the rewrite substantively outweigh the transformation overheads.
The overall benefit provided by a rewriting tool is quantified by the number of CPR obtained on the slow query workload. Additionally, we also measure CSGM, the geometric mean (GM) of the cost speedups obtained by these rewrites. Finally, to assess the actual run-time benefits, we evaluate TSGM, the geometric mean (GM) of the response-time speedups obtained by these rewrites.
@@ 2.2 Query Micro-benchmark
To motivate the progression of strategies incorporated in LITHE, we create an initial micro-benchmark comprising 10 diverse TPC-DS queries for which we were able to hand-craft high-quality CPRs. These queries are processed on GPT-4o, the popular OpenAI LLM, and the rewrites are evaluated on the PostgreSQL v16 database engine. The human rewrites deliver a CSGM of 11.84, serving as an aspirational target to attain computationally. Later, in Section 7, we extend the evaluation to complete benchmarks.
Further, for ease of presentation, we focus on the CSGM metric in the LITHE design sections (Sec 345). The TSGM performance is subsequently discussed in Section 7.
@@ 3 Basic Prompts
In this section, we explore the simplest interface to LLMs, namely prompting, for query rewriting. We evaluate four basic prompts, enumerated in Figure 4, which cover a progressive range of instructional detail and test the effectiveness of the LLM’s base knowledge.
@@
Prompt 3: Verbose instructions are given to rewrite the query, providing step-by-step guidance to the LLM to think rationally. It is first asked to pick out potential inefficiencies in the input query, and then tasked to identify approaches to address these inefficiencies. Finally, it is instructed to apply the identified solution. Essentially, the prompt tries to make the LLM reason akin to human experts.
@@
The performance of the four prompt templates on the micro-benchmark is shown in Table 1. We find that less than half the rewrites are productive with individual prompts. However, a drill-down shows that the best prompt in the ensemble is query-specific – this opens up the possibility of using all four prompts in parallel, and then choosing the best among them. This ensemble approach raises the CPRs to 6 (Row 5 in Table 1) – however, there remain four queries that are not productively rewritten by these prompts.
Table 1: Performance of Basic Prompts.
Prompt # CPR CSGM
Prompt 1 3 1.99
Prompt 2 2 1.85
Prompt 3 4 2.83
Prompt 4 4 3.00
Prompt Ensemble 6 3.23
SOTA Ensemble 3 2.49
The CSGM, shown in the last column of Table 1, is at most 3 for the individual prompts, while the ensemble reaches 3.23. But these speedups, although productive, are all lower than those delivered by the human rewrites.
Finally, an ensemble of SOTA techniques (described in Section 7) was also processed on the same platform. They delivered 3 CPRs with a CSGM of 2.49 (last row in Table 1), indicating the wide gap between the current reality and what is humanly possible.